OpenStudy (anonymous):

A student takes ten exams during a semester and receives the following grades: 90, 85, 97, 76, 89, 58, 82, 102, 70, and 67. Find the five-number summary used in a boxplot.

OpenStudy (texaschic101):

put the numbers in order 58, 67, 70, 76, 82, 85, 89, 90, 97, 102 minimum (lowest number) = 58 maximum (highest number) = 102 median (Q2) - middle number = (82 + 85) / 2 = 167/2 = 83.5 Q1 = 70 Q3 = 90